OpenStudy (anonymous):

a sharpshooter misses a target 10% of the time. what is the probability that the first miss comes after the 3rd shot? what is the probability that her 2nd miss comes no later than her 10th shot?

OpenStudy (anonymous):

let H and M be hit and miss event respectively then i)P(HHHM)=P(H)*P(H)*P(H)*P(M)=9/10*9/10*9/10*1/10=0.0729 II)P(MHHHHHHHHHM)+P(HMHHHHHHHHM)+P(HHMHHHHHHHM)+....+P(HHHHHHHHHMM) FIND THIS VALUE..AND GETB YOUR ANSWER